# Strong constraints on a sharp change in $G$ as a solution to the Hubble tension

### v2

@@ Abstract
It has been proposed that the gravitational constant $G$ abruptly decreased around 130 Myr ago, making Type Ia supernovae (SNe) in the Hubble flow intrinsically brighter than those in host galaxies with Cepheid distances. This would make Hubble flow SNe more distant, causing redshifts to rise slower with distance, potentially solving the Hubble tension. We explore a wide range of unattractive consequences of this “$G$ step model” (GSM). We find that since the luminosities of Sun-like stars scale as approximately $G^{5.6}$, the Solar luminosity would have dropped substantially 130 Myr ago in this scenario, likely pushing Earth into a planetary glaciation. However, there was no Snowball Earth episode in the last 500 Myr. The GSM also implies that the length of a year would have abruptly increased by about 10%, but the number of days per year has evolved broadly continuously according to geochronometry and cyclostratigraphy. The GSM would drastically alter stellar evolution, causing the Sun to have exhausted about 2/3 of its fuel supply rather than 1/2. This would lead make the Sun’s helioseismic age exceed that of the oldest meteorite samples, but these agree excellently in practice. The expected age of the Universe also agrees well with that of the oldest Galactic stars assuming constant $G$. The GSM however implies these stars are younger, creating a lack of stars from the first 3 Gyr of cosmic history. These arguments cast severe doubt on the viability of the GSM: the solution to the Hubble tension should be sought elsewhere.
@@
pubyear: 2025pagerange: Strong constraints on a sharp change in $G$ as a solution to the Hubble tensionReferences
@@ 1 Introduction
Cosmology is currently in a crisis because the redshift $z$ of objects in the local Universe increases with their distance $r$ more steeply than expected in the standard cosmological paradigm known as $\Lambda$-Cold Dark Matter ($\Lambda$CDM; Efstathiou, Sutherland & Maddox, 1990; Ostriker & Steinhardt, 1995) if its parameters are calibrated using the pattern of anisotropies in the cosmic microwave background (CMB). Measurements of the CMB with the Planck satellite imply that in a $\Lambda$CDM context, the Hubble constant $H_{0}=67.6\pm 0.5$ km/s/Mpc (Tristram et al., 2024), a value we denote $H_{0}^{\mathrm{Planck}}$ (Planck Collaboration VI, 2020). In a homogeneously expanding universe, this should be the local redshift gradient $cz^{\prime}$, where $c$ is the speed of light and $z^{\prime}\equiv dz/dr$. However, observations using a variety of distance indicators show that $cz^{\prime}$ is about 10% larger (Uddin et al., 2024). This discrepancy is known as the Hubble tension (for a review, see Di Valentino, 2021).
@@
Various solutions have been proposed for the Hubble tension (for a review, see Di Valentino et al., 2021). One can question whether observations of the CMB really imply such a low $H_{0}$, which is not the case in early dark energy models (Poulin et al., 2023, and references therein). If we avoid the difficulties with such approaches by assuming that $H_{0}=H_{0}^{\mathrm{Planck}}$ (Vagnozzi, 2021, 2023; Toda et al., 2024), we can alter the expansion history at late times (Rezazadeh et al., 2024) or question if $H_{0}=cz^{\prime}$, an assumption that is violated if we are living in a large local void (Keenan et al., 2013; Haslbauer et al., 2020; Mazurenko et al., 2024, 2025). Here, we focus on a possible scenario in which the local $cz^{\prime}$ is smaller than typically quoted in the literature. Since the redshifts are spectroscopically determined, such a scenario requires that the distances to the relevant objects be larger than published. The main route to measuring the local $cz^{\prime}$ involves Type Ia supernovae (SNe) in the ‘Hubble flow’ at redshifts of about $0.023-0.15$ or distances of about $100-600$ Mpc, where peculiar velocities should have little effect on $cz^{\prime}$ (Camarena & Marra, 2018, 2020a, 2020b). To calibrate the absolute SN magnitude, it is necessary to use SNe at distances $\la 40$ Mpc so that the distance to the host galaxy can be determined, typically using Cepheid variables or the tip of the red giant branch (TRGB; Baade, 1944; Li & Beaton, 2024). The period-luminosity relation of Cepheid variables (the Leavitt law; Leavitt, 1907) and the TRGB magnitude can be calibrated in the Milky Way using geometric distances.
@@
The Cepheid–SNe route is the most well-established way to obtain the local $cz^{\prime}$. However, a problem with one or more links in this ‘distance ladder’ could potentially invalidate the measurement. In particular, it is necessary to assume that the same Leavitt law continues to hold in galaxies too distant for geometric anchors. Likewise, SNe in host galaxies with Cepheid distances must be similar to the more distant SNe in the Hubble flow. The proposal we will focus on questions this last link in the logical chain by postulating that the gravitational constant $G$ changed abruptly $\ga 130$ Myr ago, which corresponds to distances $\ga 40$ Mpc, beyond the present range of Cepheid calibration. The $G$ step model (GSM; Marra & Perivolaropoulos, 2021; Perivolaropoulos & Skara, 2021; Perivolaropoulos, 2024; Ruchika et al., 2024a, b) exploits this gap between the outer limit of Cepheid distances to SN host galaxies and the inner limit to what can be considered the Hubble flow. Since the Leavitt law is calibrated empirically and the terrestrial value of $G$ applies to the entire Cepheid calibration zone, the GSM does not affect the Cepheid distances to SN host galaxies.
@@
The GSM relies on the fact that $G$ affects the SN luminosity $L$ that enters into cosmological analyses (Wright & Li, 2018; Zhao et al., 2018). The calculation is complicated somewhat by the fact that Type Ia SNe are not standard candles but are standardisable (as indeed are Cepheids). It is well known that the peak luminosity correlates with the time required for the SN light curve to decay (Phillips, 1993; Phillips et al., 1999). Changing $G$ affects both the peak luminosity and the decay time. Cosmological analyses correct SNe for the latter (and also for the SN colour) using the Tripp formula (Tripp, 1998; Brout et al., 2022). It is therefore necessary to use the Tripp-corrected SN luminosity $L$, which has the value $L_{0}$ for the terrestrially measured gravitational constant of $G_{0}$. We use Figure 1 to show the power-law fit to the Wright & Li (2018) relation between $L/L_{0}$ and $G/G_{0}$ ($L\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}G^{1.46}$; see appendix B of Desmond et al., 2019).
@@
Figure 1: Comparison of constraints on and requirements of the GSM. The black line shows the dependence of the luminosity of a Type Ia SN on $G$ in units of the terrestrial value $G_{0}$ (see figure 1 of Zhao, Wright & Li, 2018). This calculation accounts for changes in the shape of the light curve (Wright & Li, 2018), which impacts the standardised SN luminosity $L$ that enters cosmological analyses. Following appendix B of Desmond et al. (2019), we show their $L\propto G^{1.46}$ power-law fit to the results of Wright & Li (2018). The horizontal red band shows the required value of $L$ in units of the standard value $L_{0}$ that would be required to solve the Hubble tension by inflating the distances to SNe in the Hubble flow (Equation 1). In the GSM, these SNe were brighter than nearby SNe in host galaxies with Cepheid distances because $G$ was larger than today $\ga 130$ Myr ago. The vertical green band shows the cosmological constraint on $G$ using the latest CMB and baryon acoustic oscillation data (Lamine et al., 2024). It is evident that the required enhancement to $G$ is in tension with cosmological constraints, even though the aim of the GSM is to preserve the Planck cosmology. This is even before the more stringent Solar System-scale constraints described in the text.
@@
Since the GSM aims to preserve the Planck cosmology, one may worry whether a higher $G$ prior to recombination would affect the CMB anisotropies. This is an important issue given that CMB observations probe well into the diffusion damping tail, which does in fact constrain the pre-transition value of $G$ relevant to cosmology (Ooba et al., 2016, 2017). These constraints have steadily improved (Ballardini et al., 2022; Sakr & Sapone, 2022) and now limit the cosmologically relevant value of $G$ to within a few percent of the terrestrial value (Lamine et al., 2024). As the cosmological observables they consider are barely sensitive to the cosmologically brief post-transition period where $G$ had its modern value, we can interpret their study as constraining the pre-transition $G$. We show their $1\sigma$ and $2\sigma$ constraints using the vertical green bands in Figure 1. It is clear that the GSM struggles to explain early-Universe observables like the CMB consistently with the local $cz^{\prime}$ measured via the Cepheid–SNe route.
In this work, we consider several other major problems with the GSM on a variety of scales. Section 2 considers how the proposed sharp drop in $G$ would affect the Sun and the rest of the Solar System, especially the Earth. In Section 3, we consider the impact on the evolution of other stars and how that would affect other distance indicators and measures of cosmic expansion. We summarise our results in Section 4 and conclude by considering empirically viable model extensions. In Appendix A, we consider a few more subtle consequences of the GSM and discuss how these might be constrained by future investigations.
@@ 2 Impact on the Solar System
The GSM has drastic consequences in the Solar System, which we discuss in this section. The effects on the Sun (Section 2.1) would in turn affect Earth, both in terms of its climate (Section 2.2) and its orbit (Section 2.3). On a smaller scale, a sharp drop in $G$ would also affect the Earth–Moon system (Section 2.4).
@@ 2.2 Impact on the terrestrial climate
The GSM implies a severe drop in $L_{\odot}$, but the sharp drop in $G$ would also expand the Earth’s orbital radius $r$. We will return to this issue in Section 2.3, but our main concern here is that this leads to a further drop in the Solar insolation on the Earth. Taking into account that Earth’s specific angular momentum is $\sqrt{GM_{\odot}r}$ given its nearly circular orbit, a drop in $G$ over several years increases $r$ such that $r\propto 1/G$ in order to conserve angular momentum. Combined with the fact that $L_{\odot}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}G^{5.6}$, the Earth’s blackbody temperature $T_{\oplus}\propto G^{1.9}$. [^1] This is somewhat shallower than the earlier estimate that $T_{\oplus}\propto G^{9/4}$ (Teller, 1948) because that study assumed $L_{\odot}\propto G^{7}$.
@@
In reality, the Earth is of course not a perfect blackbody in equilibrium with the incident Solar radiation. On long timescales, its climate is maintained in the temperature range that allows liquid water oceans thanks to the carbonate-silicate cycle, which relies on the temperature dependence of weathering processes that remove the greenhouse gas CO2 (Walker, Hays & Kasting, 1981). However, reduced weathering would take many Myr to allow volcanic outgassing to build up enough atmospheric CO2 to compensate for such a drastic reduction in Solar insolation. It is therefore inevitable that the Earth would experience a period of significantly reduced temperature. Given that the global average $T_{\oplus}\approx 290$ K today, a $>5\%$ drop in $G$ would have reduced $T_{\oplus}$ by $\ga 10\%$ to temperatures below the freezing point of water. As water turned to ice near the polar regions, the surface would have become more reflective, further reducing the temperature. This ice-albedo feedback effect (Budyko, 1969; Sellers, 1973) would most likely lead to a planetary glaciation lasting many Myr.
@@
Historically, there is actually very strong evidence for several such Snowball Earth episodes (for a review, see Banik, 2016). However, these all took place $\ga 600$ Myr ago. This corresponds to a distance of almost 200 Mpc, where we can assume a fairly smooth distance-redshift relation. It is not possible to put the GSM that far back in the past because this would lead to a severe mismatch between the luminosities of Hubble flow SNe with redshifts either side of the transition, which have a similar distance. A feature would be produced in the SN Hubble diagram, which is not observed. The only way to ‘hide’ a substantial transition in $L_{0}$ (and hence in apparent magnitude at given $z$) is if it occurred sufficiently nearby that the redshift cannot be used to reliably determine the distance and thus lookback time. Peculiar velocities would ‘blur out’ such a local transition, but not one in the Hubble flow region that starts $\ga 100$ Mpc away in standard cosmology. The transition must therefore lie closer to us, and hence cannot be matched to an observed Snowball Earth episode (see Section 4 for the possibility that the issue is instead in the lowest rung of the distance ladder). The lack of any Snowball Earth episode in the past 500 Myr implies that the proposed sharp drop in $G$ did not occur within 150 Mpc, which however is necessary for the GSM to solve the Hubble tension. This already casts very severe doubt on the GSM. An important caveat is that the response of the terrestrial climate to a sharp drop in $L_{\odot}$ needs to be explored in more detail to assess if a planetary glaciation could have been avoided, though models suggest that even a modest reduction in $L_{\odot}$ could trigger a planetary glaciation (Arnscheidt & Rothman, 2020). If however this can be avoided, then a more modest drop in $T_{\oplus}$ might be hidden in the geological record given other causes of temperature fluctuations and measurement uncertainties.

### v4

# Challenges to a sharp change in $G$ as a solution to the Hubble tension

@@ Abstract
It has been proposed that the gravitational constant $G$ abruptly decreased around 130 Myr ago, making Type Ia supernovae (SNe) in the Hubble flow intrinsically brighter than those in host galaxies with Cepheid distances. This would make Hubble flow SNe more distant, causing redshifts to rise slower with distance, potentially solving the Hubble tension. We explore a wide range of unattractive consequences of this “$G$ step model” (GSM). We find that since the luminosities of Sun-like stars scale as approximately $G^{5.6}$, the Solar luminosity would have dropped substantially 130 Myr ago in this scenario, likely pushing Earth into a planetary glaciation. However, there was no Snowball Earth episode in the last 500 Myr. The GSM also implies that the length of a year would have abruptly increased by about 10%, but the number of days per year has evolved broadly continuously according to geochronometry and cyclostratigraphy. The GSM would considerably alter stellar evolution, causing the Sun to have exhausted about 2/3 of its fuel supply rather than 1/2. This would make the Sun’s helioseismic age exceed that of the oldest meteorite samples, but these agree excellently in practice. The expected age of the Universe also agrees well with that of the oldest Galactic stars assuming constant $G$. The GSM however implies these stars are younger, creating a lack of stars from the first 3 Gyr of cosmic history. These arguments pose significant challenges to models seeking to resolve the Hubble tension through a transition in $G$.
@@
pubyear: 2025pagerange: Challenges to a sharp change in $G$ as a solution to the Hubble tensionA
@@ 1 Introduction
Cosmology is currently in a crisis because the redshift $z$ of objects in the local Universe increases with their distance $r$ more steeply than expected in the standard cosmological paradigm known as $\Lambda$-Cold Dark Matter ($\Lambda$CDM; Efstathiou, Sutherland & Maddox, 1990; Ostriker & Steinhardt, 1995) if its parameters are calibrated using the pattern of anisotropies in the cosmic microwave background (CMB). Measurements of the CMB with the Planck satellite imply that in a $\Lambda$CDM context, the Hubble constant $H_{0}=67.6\pm 0.5$ km/s/Mpc (Tristram et al., 2024), a value we denote $H_{0}^{\mathrm{Planck}}$ (Planck Collaboration VI, 2020). In a homogeneously expanding universe, this should be the local redshift gradient $cz^{\prime}$, where $c$ is the speed of light and $z^{\prime}\equiv dz/dr$. However, observations using a variety of distance indicators show that $cz^{\prime}$ is about 10% larger (Riess et al., 2024; Uddin et al., 2024; Scolnic et al., 2025). This discrepancy is known as the Hubble tension (for a review, see Di Valentino, 2021).
@@
Various solutions have been proposed for the Hubble tension (for a review, see Di Valentino et al., 2021). One can question whether observations of the CMB really imply such a low $H_{0}$, which is not the case in early dark energy models (Poulin et al., 2023, and references therein). If we avoid the difficulties with such approaches by assuming that $H_{0}=H_{0}^{\mathrm{Planck}}$ (Vagnozzi, 2021, 2023; Toda et al., 2024) and trust the local $cz^{\prime}$, we would have to question if $H_{0}=cz^{\prime}$, an assumption that is violated if we are living in a large local void (Keenan et al., 2013; Haslbauer et al., 2020; Mazurenko et al., 2024, 2025; Banik & Kalaitzidis, 2025). We could instead alter the expansion history at late times (Rezazadeh et al., 2024). Here, we focus on a possible scenario in which $H_{0}=cz^{\prime}$ but the local $cz^{\prime}$ is smaller than typically quoted in the literature. Since the redshifts are spectroscopically determined, such a scenario requires that the distances to the relevant objects be larger than published. The main route to measuring the local $cz^{\prime}$ involves Type Ia supernovae (SNe) in the ‘Hubble flow’ at redshifts of about $0.023-0.15$ or distances of about $100-600$ Mpc, where peculiar velocities should have little effect on $cz^{\prime}$ (Camarena & Marra, 2018, 2020a, 2020b). To calibrate the absolute SN magnitude, it is necessary to use SNe at distances $\la 40$ Mpc so that the distance to the host galaxy can be determined, typically using Cepheid variables or the tip of the red giant branch (TRGB; Baade, 1944; Li & Beaton, 2024). The period-luminosity relation of Cepheid variables (the Leavitt law; Leavitt, 1907) and the TRGB magnitude can be calibrated in the Milky Way using geometric distances.
@@
The Cepheid–SNe route is the most well-established method to obtain the local $cz^{\prime}$. However, a problem with one or more links in this ‘distance ladder’ could potentially invalidate the measurement. In particular, it is necessary to assume that the same Leavitt law continues to hold in galaxies too distant for geometric anchors. Likewise, SNe in host galaxies with Cepheid distances must be similar to the more distant SNe in the Hubble flow. The proposal we will focus on questions this last link in the logical chain by postulating that the gravitational constant $G$ changed abruptly $\ga 130$ Myr ago, which corresponds to distances $\ga 40$ Mpc, beyond the present range of Cepheid calibration. The $G$ step model (GSM; Marra & Perivolaropoulos, 2021; Perivolaropoulos & Skara, 2021; Perivolaropoulos, 2024; Ruchika et al., 2024a, b) exploits this gap between the outer limit of Cepheid distances to SN host galaxies and the inner limit to what can be considered the Hubble flow. Since the Leavitt law is calibrated empirically and the terrestrial value of $G$ applies to the entire Cepheid calibration zone, the GSM does not affect the Cepheid distances to SN host galaxies.
@@
The GSM relies on the fact that $G$ affects the SN luminosity $L$ that enters into cosmological analyses (Wright & Li, 2018; Zhao et al., 2018). The calculation is complicated somewhat by the fact that Type Ia SNe are not standard candles but are standardizable (as indeed are Cepheids). It is well known that the peak luminosity correlates with the time required for the SN light curve to decay (Phillips, 1993; Phillips et al., 1999). Changing $G$ affects both the peak luminosity and the decay time. Cosmological analyses correct SNe for the latter (and also for the SN colour) using the Tripp formula (Tripp, 1998; Brout et al., 2022). It is therefore necessary to use the Tripp-corrected SN luminosity $L$, which has the value $L_{0}$ for the terrestrially measured gravitational constant of $G_{0}$. We use Figure 1 to show the power-law fit to the Wright & Li (2018) relation between $L/L_{0}$ and $G/G_{0}$ ($L\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}G^{1.46}$; see appendix B of Desmond et al., 2019).
@@
Figure 1: Comparison of constraints on and requirements of the GSM. The black line shows the dependence of the luminosity of a Type Ia SN on $G$ in units of the terrestrial value $G_{0}$ (see figure 1 of Zhao, Wright & Li, 2018). This calculation accounts for changes in the shape of the light curve (Wright & Li, 2018), which impacts the standardized SN luminosity $L$ that enters cosmological analyses. Following appendix B of Desmond et al. (2019), we show their $L\propto G^{1.46}$ power-law fit to the results of Wright & Li (2018). The horizontal red band shows the required value of $L$ in units of the standard value $L_{0}$ that would be needed to solve the Hubble tension by inflating the distances to SNe in the Hubble flow (Equation 1). In the GSM, these SNe were brighter than nearby SNe in host galaxies with Cepheid distances because $G$ was larger than today $\ga 130$ Myr ago. The vertical green band shows the cosmological constraint on $G$ using the latest CMB and baryon acoustic oscillation data (Lamine et al., 2025). It is evident that the required enhancement to $G$ is in tension with cosmological constraints, even though the aim of the GSM is to preserve the Planck cosmology. This is even before the more stringent Solar System-scale constraints described in the text.
@@
Since the GSM aims to preserve the Planck cosmology, one may worry whether a higher $G$ prior to recombination would affect the CMB anisotropies. This is an important issue given that CMB observations probe well into the diffusion damping tail, which does in fact constrain the pre-transition value of $G$ relevant to cosmology (Ooba et al., 2016, 2017). These constraints have steadily improved (Ballardini et al., 2022; Sakr & Sapone, 2022) and now limit the cosmologically relevant value of $G$ to within a few percent of the terrestrial value (Lamine et al., 2025). As the cosmological observables they consider are barely sensitive to the cosmologically brief post-transition period where $G$ had its modern value, we can interpret their study as constraining the pre-transition $G$. We show their $1\sigma$ and $2\sigma$ constraints using the vertical green bands in Figure 1. It is clear that the GSM struggles to explain early-Universe observables like the CMB consistently with the local $cz^{\prime}$ measured via the Cepheid–SNe route.
In this work, we consider several other significant challenges to building an observationally viable GSM. Section 2 considers how the proposed sharp drop in $G$ would affect the Sun and the rest of the Solar System, especially the Earth. In Section 3, we consider the impact on the evolution of other stars and how that would affect other distance indicators and measures of cosmic expansion. We summarize our results in Section 4 and conclude by considering empirically viable model extensions. In Appendix A, we consider a few more subtle consequences of the GSM and discuss how these might be constrained by future investigations.
@@ 2 Impact on the Solar System
The GSM has important detrimental consequences in the Solar System, which we discuss in this section. The effects on the Sun (Section 2.1) would in turn affect Earth, both in terms of its climate (Section 2.2) and its orbit (Section 2.3). On a smaller scale, a sharp drop in $G$ would also affect the Earth–Moon system (Section 2.4).
@@ 2.2 Impact on the terrestrial climate
The GSM implies a significant drop in $L_{\odot}$, but the sharp drop in $G$ would also expand the Earth’s orbital radius $r$. We will return to this issue in Section 2.3, but our main concern here is that this leads to a further drop in the Solar insolation on the Earth. Taking into account that Earth’s specific angular momentum is $\sqrt{GM_{\odot}r}$ given its nearly circular orbit, a drop in $G$ over several years increases $r$ such that $r\propto 1/G$ in order to conserve angular momentum. Combined with the fact that $L_{\odot}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}G^{5.6}$, the Earth’s blackbody temperature $T_{\oplus}\propto G^{1.9}$. [^1] This is somewhat shallower than the earlier estimate that $T_{\oplus}\propto G^{9/4}$ (Teller, 1948) because that study assumed $L_{\odot}\propto G^{7}$.
@@
In reality, the Earth is of course not a perfect blackbody in equilibrium with the incident Solar radiation. On long timescales, its climate is maintained in the temperature range that allows liquid water oceans thanks to the carbonate-silicate cycle, which relies on the temperature dependence of weathering processes that remove the greenhouse gas CO2 (Walker, Hays & Kasting, 1981). However, reduced weathering would take many Myr to allow volcanic outgassing to build up enough atmospheric CO2 to compensate for such a sizeable reduction in Solar insolation. It is therefore inevitable that the Earth would experience a period of significantly reduced temperature. Given that the global average $T_{\oplus}\approx 290$ K today, a $>5\%$ drop in $G$ would have reduced $T_{\oplus}$ by $\ga 10\%$ to temperatures below the freezing point of water. As water turned to ice near the polar regions, the surface would have become more reflective, further reducing the temperature. This ice-albedo feedback effect (Budyko, 1969; Sellers, 1973) would most likely lead to a planetary glaciation lasting many Myr.
@@
Historically, there is actually very strong evidence for several such Snowball Earth episodes (for a review, see Banik, 2016). However, these all took place $\ga 600$ Myr ago. This corresponds to a distance of almost 200 Mpc, where we can assume a fairly smooth distance-redshift relation. It is not possible to put the $G$ transition that far back in the past because this would lead to a clear mismatch between the luminosities of Hubble flow SNe with redshifts either side of the transition, which have a similar distance. A feature would be produced in the SN Hubble diagram, which is not observed. The only way to ‘hide’ a substantial transition in $L_{0}$ (and hence in apparent magnitude at given $z$) is if it occurred sufficiently nearby that the redshift cannot be used to reliably determine the distance and thus lookback time. Peculiar velocities would ‘blur out’ such a local transition, but not one in the Hubble flow region that starts $\ga 100$ Mpc away in standard cosmology. The transition must therefore lie closer to us, and hence cannot be matched to an observed Snowball Earth episode (see Section 4 for the possibility that the issue is instead in the lowest rung of the distance ladder). The lack of any Snowball Earth episode in the past 500 Myr implies that the proposed sharp drop in $G$ did not occur within 150 Mpc, which however is necessary for the GSM to solve the Hubble tension. This already casts significant doubt on the GSM.
An important caveat is that the response of the terrestrial climate to a sharp drop in $L_{\odot}$ needs to be explored in more detail to assess if a planetary glaciation could have been avoided, though models suggest that even a modest reduction in $L_{\odot}$ could trigger a planetary glaciation (Arnscheidt & Rothman, 2020). If however this can be avoided, then a more modest drop in $T_{\oplus}$ might be hidden in the geological record given other causes of temperature fluctuations and measurement uncertainties.